OpenStudy (dmezzullo):

The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected!

OpenStudy (anonymous):

guess and check for this one

OpenStudy (anonymous):

x + (x+1) +(x+2) = 267 solve

OpenStudy (anonymous):

yeh follow @freewilly922 's method.

OpenStudy (anonymous):

divide 267 by 3. the answer is the middle integer, the number before and the number after are the others.

OpenStudy (dmezzullo):

u lost me

OpenStudy (dmezzullo):

easiest terms possible plz

OpenStudy (aripotta):

x + (x+1) +(x+2) = 267. x being the first number, (x + 1) being the next consecutive number, and (x + 2) being the second consecutive number. the sum of those equal 267

OpenStudy (aripotta):

so x + x + x = 3x and 1 + 2 = 3. so 3x + 3 = 267

OpenStudy (dmezzullo):

tht the answ?

OpenStudy (aripotta):

no, you gotta solve for x first

OpenStudy (aripotta):

then you can solve (x + 2)

OpenStudy (aripotta):

which would be the largest number, right?

OpenStudy (dmezzullo):

me=lost gain

OpenStudy (aripotta):

:/

OpenStudy (anonymous):

Which is why he should guess the numbers. (Guess and check like satellite prescribed)

OpenStudy (aripotta):

no, he shouldn't guess. he already knows 3x + 3 = 267, all he has to do is solve for x, and add 2..

OpenStudy (dmezzullo):

91?

OpenStudy (aripotta):

it's 90

OpenStudy (aripotta):

x = 88, 88 + 2 = 90

OpenStudy (aripotta):

does that make sense?

OpenStudy (dmezzullo):

yea buddie

OpenStudy (dmezzullo):

oh so 88,90 r the answers?

OpenStudy (aripotta):

no, 90 is the answer. it only asked for the largest integer

OpenStudy (aripotta):

88 is the first integer. 89 is the next number, and 90 is the largest

OpenStudy (dmezzullo):

k

OpenStudy (aripotta):

88 + 89 + 90 = 267

OpenStudy (dmezzullo):

ohhhh lol sorry my brain aint funtioning right its to late

OpenStudy (anonymous):

User: Taylor Bennett In Course: Algebra I V17 ( 3939) Instructor: Jessica Smith 02.02 Two-Variable Equations Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. -------------------------------------------------------------------------------- Question 1 (Fill-In-The-Blank Worth 1 points) (02.02) A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol. Numerical Answers Expected! Answer for Blank 1: -------------------------------------------------------------------------------- Question 2 (Fill-In-The-Blank Worth 2 points) (02.02) The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds Numerical Answers Expected! Answer for Blank 1: -------------------------------------------------------------------------------- Question 3 (Fill-In-The-Blank Worth 1 points) (02.02) Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values. Numerical Answers Expected! Answer for Blank 1: -------------------------------------------------------------------------------- Question 4 (Fill-In-The-Blank Worth 3 points) (02.02) Evan is making a table that will be created in the shape of the figure below. The table top is a triangle attached to a rectangle. To purchase the right amount of paint, he needs to know the area of the table top. He can only spend $10 on paint, which is enough to cover 150 ft2 of surface area. What is the maximum length of the base of the rectangle he can build? x = __________________ ft Numerical Answers Expected! Answer for Blank 1: -------------------------------------------------------------------------------- Question 5 (Fill-In-The-Blank Worth 1 points) (02.02) The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected! Answer for Blank 1: -------------------------------------------------------------------------------- Question 6 (Fill-In-The-Blank Worth 2 points) (02.02) As the number of pages in a photo book increases, the price of the book also increases. There is an additional shipping charge of 15%. The price of a book can be modeled by the equation below where, P = the price of the book, 20 is the printing charge, 0.5 is the charge per page, and x = the number of pages. P = (20 + 0.5x) + 0.15(20 + 0.5x) Jennifer wants to purchase a book but only has $62.10 to spend. What is the maximum number of pages she can have in her book?